# Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents

### v3

@@ Abstract
Recent advancements on Large Language Models (LLMs) enable AI Agents to automatically generate and execute multi-step plans to solve complex tasks. However, since LLM’s content generation process is hardly controllable, current LLM-based agents frequently generate invalid or non-executable plans, which jeopardizes the performance of the generated plans and corrupts users’ trust in LLM-based agents. In response, this paper proposes a novel “Formal-LLM” framework for LLM-based agents by integrating the expressiveness of natural language and the precision of formal language. Specifically, the framework allows human users to express their requirements or constraints for the planning process as an automaton. A stack-based LLM plan generation process is then conducted under the supervision of the automaton to ensure that the generated plan satisfies the constraints, making the planning process controllable. We conduct experiments on both benchmark tasks and practical real-life tasks, and our framework achieves over 50% overall performance increase, which validates the feasibility and effectiveness of employing Formal-LLM to guide the plan generation of agents, preventing the agents from generating invalid and unsuccessful plans. Further, more controllable LLM-based agents can facilitate the broader utilization of LLM in application scenarios where high validity of planning is essential. The work is open-sourced at https://github.com/agiresearch/Formal-LLM.
Large Language Model, Agents, Controllability
1Department of Computer Science, Rutgers University, New Brunswick, NJ 08854, USA
zelong.li, wenyue.hua, hw488, hz375, yongfeng.zhang@rutgers.edu
@@ 1 Introduction
Numerous applications have emerged with the rapid development of Large Language Models (LLM). One notable application is the LLM-based agent, which is capable of automatically generating and executing multi-step plans to solve complex tasks. While the LLM-based agent exhibits creativity, there is a concern about the potential generation of unreasonable and invalid plans, undermining the effectiveness of agents. For example, generating a plan that attempts to process image data using a tool designed for text can lead to errors. Recent studies have pointed out the challenges of LLM-based agents in developing non-executable plans without sufficient human oversight (Ge et al., 2023a; Yuan et al., 2024). Addressing these challenges is crucial for improving agent performance, increasing the generation of valid plans, and maintaining user trust. Various attempts have been made to control LLM text generation, such as incorporating hard constraints (Takase and Okazaki, 2019; Carlsson et al., 2022), soft constraints (Gu et al., 2022; Lu et al., 2022), or a combination of both (Chen et al., 2024). However, the focus of controlling LLM-based agents emphasizes the validity of plans and tool-use over purely text generation. Some recent studies leverage LLM as a parser to extract a chain of tools from the generated texts based on pre-designed prompts (Ge et al., 2023a; Yuan et al., 2024), yet achieving a satisfactory rate of valid plans remains elusive.
@@
To tackle the problems of valid plan generation, we propose a framework named “Formal-LLM”, which integrates the expressiveness of natural language and the precision of formal language, as shown by the toy example in Figure 1. Specifically, to control the LLM-based agent’s plan generation, human users construct a context-free grammar (CFG) as the formal language to represent the constraints for the agent. The CFG is then automatically translated into a pushdown automaton (PDA). When LLM conducts planning, it is prompted to follow the state transition defined by the automaton. This is realized by limiting the LLM-based agent’s choices at each step to the valid actions defined by the PDA at its current state, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA in this work because some tasks need to be solved by a tree-structured plan rather than a simple chain-structured plan, as shown by the examples in Figure 5, and such tree-structured plans require a PDA to generate.
@@
Figure 1: The Formal-LLM workflow with a toy example. To control the LLM-based agent’s plan generation, human users construct a formal language (context-free grammar, CFG) to represent the natural language constraints. The formal language is then translated into a pushdown automaton (PDA). When LLM conducts planning, it needs to follow the state transition defined by the automaton, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA because some plans exhibit tree-structure (as shown by other examples in the paper), which requires a PDA to generate.
@@ 2.1 LLM-based AI Agent
An AI agent is an autonomous entity capable of making decisions and executing actions within a specific environment to effectively tackle diverse complex tasks using AI techniques (Ge et al., 2023b; Wang et al., 2023; Xi et al., 2023). The emergence of Large Language Models (LLMs), exemplified by the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) and the LLaMA series (Touvron et al., 2023a, b), has spurred the exploration of LLM-based agents (Ge et al., 2023a; Huang et al., 2022). These agents utilize LLMs as their central cognitive component or controller, broadening their perceptual and action capabilities through approaches like multimodal perception and tool usage (Schick et al., 2023; Ge et al., 2023a; Qin et al., 2023). In contrast to pre-LLM era AI agents, LLM-based agents showcase creativity, manifesting in the ability to generate innovative ideas without additional learning (Franceschelli and Musolesi, 2023), indicating a degree of self-directed exploration and decision-making (Xi et al., 2023). Despite LLM-based agent’s application in diverse real-world scenarios such as software development (Li et al., 2023a; Qian et al., 2023), scientific research (Boiko et al., 2023), and system management (Liu et al., 2023), recent studies have highlighted issues of non-executable plans being generated without adequate human oversight (Ge et al., 2023a; Yuan et al., 2024). If these plans lack executability, the utility of LLM-based agents in fields that require high validity is compromised, and their unreliability erodes user trust. To address this challenge, we propose integrating natural language and precise automaton during the agent’s planning process to enable the controllable generation of valid plans.
@@ 2.2 Controllable LLM Generation
To the best of our knowledge, there are few works on controllable LLM-based agents, and discussions primarily revolve around controllable LLM text generation. Controlling LLM text generation typically falls into two categories: hard constraints and soft constraints (Qin et al., 2022). Hard constraints involve constraints on the desired text length, designated keywords for the generated text (Takase and Okazaki, 2019; Carlsson et al., 2022), and constraint vocabulary space in decoding time (Hemmer et al., 2023; Geng et al., 2024, 2023), while soft constraints restrict the output based on specific semantics, such as sentiments or topics (Gu et al., 2022; Lu et al., 2022; Li et al., 2020). A recent work (Chen et al., 2024) attempts to integrate both constraints into a unified approach. However, text generation control has disadvantages and problems. First, text generation control could hardly be directly applied to closed-source LLMs due to their need to access the decoding probability (Geng et al., 2024). Second, controllable LLM-based agents prioritize the validity of plan and tool use over purely text generation. Recent studies on LLM-based agents have noted limited effectiveness when employing text constraint generation on LLM-based agents and have explored using LLMs as a parser to extract planning information from generated texts (Ge et al., 2023a; Yuan et al., 2024); but relying heavily on the parser LLM’s effectiveness may not yield a satisfactory plan valid rate, especially in complex tasks. We propose an automaton-guided LLM-based agent planning to ensure that the generated plans adhere 100% to the constraints.
@@ 3.1 Context-free Grammar
There are many LLM-based agent operating tools that take multiple inputs and produce one output. For example, the tool “Visual Question Answering” receives an image and a text as the inputs and generates a text as the output. The planning involving such tools forms a tree structure, and such tree-structured plans require Context-Free Grammar (CFG) to express the rules.
The CFG consists of four components: terminals (symbols unable to be replaced), nonterminals (symbols subject to replacement), start symbol (a unique nonterminal, usually denoted as $S$), and productions (rules governing symbol substitution). The CFG production follows the format:
@@
where $A$ represents a single nonterminal, and $\alpha$ is a string comprising any combination of terminals and nonterminals. $A$ can be replaced by $\alpha$ in any situations. Eq.(2) presents a CFG example with two productions:
$\begin{split}&S\rightarrow\varepsilon\\ &S\rightarrow aSb\end{split}$ (2)
where $a,b$ are terminals, $S$ is the start symbol and the only nonterminal, and $\varepsilon$ stands for the empty string. From the start symbol, we can use these two productions to generate infinite strings, such as the empty string, $ab$, and $aabb$. Take the derivation of $aabb$ as an example. We first apply the second production twice ($S\rightarrow aSb\rightarrow aaSbb$) and then the first production once ($aaSbb\rightarrow aabb)$. The formal language, Context-Free Language (CFL), contains the set of strings of terminals, which can be derived from $S$ in a certain amount of CFG production steps.
Figure 2: A Pushdown Automaton (PDA) example. The edge represents the transition function. For example, the edge $(a,Z;SZ)$ from $q_{0}$ to $q_{0}$ means if the stack top is $Z$ and the next string letter is $a$, then the PDA remains in $q_{0}$ state and the stack top changes from $Z$ to $SZ$.
@@ 3.2 Pushdown Automaton
We use Figure 2 as an example to show how PDA works. This automaton includes a stack alphabet ($\{S,Z\}$) with an initial stack symbol $Z$, a start state $q_{0}$, an accept state set $\{q_{2}\}$, and a transit function over states. The transit functions are in the form of triplets. For example, the edge $(a,Z;SZ)$ from $q_{0}$ to $q_{0}$ means that if the stack top is $Z$ and the next letter of the string is $a$, the PDA remains in $q_{0}$ state and the stack top changes from $Z$ to $SZ$ ($S$ is the stack top now).
@@ 4.1 Motivation and Challenge
Natural language is easily understandable by humans, but it may also lack precision in certain application scenarios. In contrast, formal language is defined in clear mathematical and machine readable forms, which could be hard to comprehend for human users, but it also possesses great precision. Thus, we aim to effectively integrate the benefits of both natural and formal languages through the pushdown automaton (PDA) for more controllable planning of LLM-based agents in scenarios requiring precise and valid plans.
Constructing our Formal-LLM framework presents non-trivial challenges. The first challenge lies in the potential struggle of the LLM to directly comprehend and process formal language due to the limited CFG- or PDA-related corpus in pre-training data. Hence, natural language prompts are necessary to effectively describe the status of the automaton, bridging the gap between formal and natural language for LLM’s planning. Given that the automaton has the potential to generate an infinite number of plans, the second challenge arises because the automaton only ensures the validity and executability of the plans by adhering to defined constraints, but cannot guarantee the performance and optimality of the generated plan. Therefore, leveraging the solid natural language understanding ability of LLM and other methods, such as fine-tuning and backtracking mechanisms, becomes crucial for enhancing the planning performance.
@@ 4.2 Formulating Constraint to Automaton
If the plans for the benchmark tasks generated by the LLM-based agent deviate from the expected data format, the execution performance degrades significantly or even results in errors. For instance, consider the Colorization tool, which operates on image data as input and output. Given that image data is represented as a 3-D tensor, providing a 1-D tensor text string as input to Colorization could lead to errors, specifically due to incorrect data dimensions. Consequently, the entire plan becomes invalid and cannot be executed. In such cases, natural language may struggle to articulate these constraints precisely and concisely, whereas formal language can express them straightforwardly.
@@
We use symbols to represent different elements: $T$ for text and $I$ for image, capital letters $A$ to $F$ for the six types of tools, and lowercase letters such as $a_{i}\in A$ or $f_{j}\in F$ for specific tools, in the order of Table 3. The context-free grammar (CFG) below outlines the constraints on data formats with Polish prefix notation.
$\begin{split}&I\rightarrow AI|CT\\ &T\rightarrow BI|DT|EIT|FTT\\ &A\rightarrow a_{1}|a_{2}|a_{3}|a_{4}\\ &B\rightarrow b_{1}|b_{2}|b_{3}\\ &C\rightarrow c_{1}\\ &D\rightarrow d_{1}|d_{2}|d_{3}|d_{4}|d_{5}\\ &E\rightarrow e_{1}\\ &F\rightarrow f_{1}\end{split}$ (3)
@@
where $I\rightarrow AI|CT$ means $I$ can be replaced by $AI$ or $CT$, because $A$ is an image-to-image tool and $C$ is a text-to-image tool, as a result, when $A$ is applied on an image ($AI$) or $C$ is applied on a text ($CT$), the result is still an image. Considering a specific task, such as “Given blurry gray-scale images, how to return the object names in English step by step?”, we know that the task’s inputs are images and the final outputs should be texts. To formalize this, we introduce two additional constraints to Eq.(3):
$\begin{split}&S\rightarrow T\\ &I\rightarrow i\end{split}$ (4)
@@
i.e., the CFG starts with the symbol $T$, implying the final text format, and uses the lowercase letter $i$ to represent the input image. The combination of Eq.(3) and Eq.(4) renders the data format constraints of the specific task. For instance, $e_{1}a_{1}ib_{1}i$ is an accepted word in this formal language, representing a valid plan: “Utilize Image Classification ($b_{1}$) on the input image to obtain text and apply Colorization ($a_{1}$) on the input image to generate the output image. Then, employ Visual Question Answering ($e_{1}$) with the text and output image to derive the final text.” As mentioned in Section 3.2, the CFG can be equivalently converted into a PDA, illustrated in Figure 3. These steps allow us to transform the constraints into automata for subsequent utilization.
@@
In certain scenarios, creating an automaton can be more straightforward than formulating a formal language. Section B.4.1 describes a real-life task for daily work planning. The constraints in this case revolve around time, allowing us to represent time as distinct states, as depicted in Figure 4. The automaton initiates the design of an activity that concludes at 20:00, and a valid plan is generated when the automaton reaches 10:00 while being in a state where three meals have been consumed. This approach involves designing an effective PDA without the need for a CFG.

### v4

@@ Abstract
Recent advancements on Large Language Models (LLMs) enable AI Agents to automatically generate and execute multi-step plans to solve complex tasks. However, since LLM’s content generation process is hardly controllable, current LLM-based agents frequently generate invalid or non-executable plans, which jeopardizes the performance of the generated plans and corrupts users’ trust in LLM-based agents. In response, this paper proposes a novel “Formal-LLM” framework for LLM-based agents by integrating the expressiveness of natural language and the precision of formal language. Specifically, the framework allows agent developers to express their requirements or constraints for the planning process as an automaton. A stack-based LLM plan generation process is then conducted under the supervision of the automaton to ensure that the generated plan satisfies the constraints, making the planning process controllable. We conduct experiments on both benchmark tasks and practical real-life tasks, and our framework achieves over 50% overall performance increase, which validates the feasibility and effectiveness of employing Formal-LLM to guide the plan generation of agents, preventing the agents from generating invalid and unsuccessful plans. Further, more controllable LLM-based agents can facilitate the broader utilization of LLM in application scenarios where high validity of planning is essential. The source code of this work is available at https://github.com/agiresearch/Formal-LLM.
Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents
Zelong Li, Wenyue Hua, Hao Wang, He Zhu, Yongfeng Zhang Department of Computer Science, Rutgers University, New Brunswick {zelong.li, wenyue.hua, hw488, hz375, yongfeng.zhang}@rutgers.edu
@@ 1 Introduction
Numerous applications have emerged with the rapid development of Large Language Models (LLM). One notable application is the LLM-based agent, which is capable of automatically generating and executing multi-step plans to solve complex tasks. While the LLM-based agent exhibits creativity, there is a concern about the potential generation of unreasonable and invalid plans, undermining the effectiveness of agents. For example, generating a plan that attempts to process image data using a tool designed for text can lead to errors. Recent studies have pointed out the challenges of LLM-based agents in developing non-executable plans without sufficient human oversight (Ge et al., 2023a; Yuan et al., 2024). Addressing these challenges is crucial for improving agent performance, increasing the generation of valid plans, and maintaining user trust. Various attempts have been made to control LLM text generation, such as incorporating hard constraints (Takase and Okazaki, 2019; Carlsson et al., 2022), soft constraints (Gu et al., 2022; Lu et al., 2022), or a combination of both (Chen et al., 2024). However, the focus of controlling LLM-based agents emphasizes the validity of plans and tool use over purely text generation. Some studies (Ge et al., 2023a; Yuan et al., 2024) use LLM as a parser to extract a chain of tools from the generated texts based on prompts, yet satisfactory ratios for valid plans remain elusive.
@@
To tackle the problems of invalid plan generation, we propose a framework named “Formal-LLM”, which integrates the expressiveness of natural language and the precision of formal language, as shown by the toy example in Figure 1. Specifically, to control the LLM-based agent’s plan generation, agent developers construct a context-free grammar (CFG) as the formal language to represent the constraints for the agent. The CFG is then automatically translated into a pushdown automaton (PDA). When LLM conducts planning, it is prompted to follow the state transition defined by the automaton. This is realized by limiting the LLM-based agent’s choices at each step to the valid actions defined by the PDA at its current state, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA in this work because some tasks need to be solved by tree-structured plans rather than chain-structured plans, and generating tree-structured plans requires a PDA.
@@
Figure 1: The Formal-LLM workflow with a toy example. To control the LLM-based agent’s plan generation, agent developers construct a formal language (context-free grammar, CFG) to represent the natural language constraints. The formal language is then translated into a pushdown automaton (PDA). When LLM conducts planning, it needs to follow the state transition defined by the automaton, which helps to guarantee that the constraint is satisfied in the final generated plan. We choose PDA because some plans exhibit tree-structure (as shown by other examples in the paper), which requires a PDA to generate.
@@ 2.1 LLM-based AI Agent
An AI agent is an autonomous entity capable of making decisions and executing actions within a specific environment to effectively tackle diverse complex tasks using AI techniques (Ge et al., 2023b; Mei et al., 2024; Wang et al., 2023; Xi et al., 2023). The emergence of Large Language Models (LLMs), exemplified by the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) and the LLaMA series (Touvron et al., 2023a, b), has spurred the exploration of LLM-based agents (Ge et al., 2023a; Huang et al., 2022). These agents utilize LLMs as their central cognitive component or controller, broadening their perceptual and action capabilities through approaches like multimodal perception and tool usage (Schick et al., 2023; Ge et al., 2023a; Qin et al., 2023). In contrast to pre-LLM era AI agents, LLM-based agents showcase creativity, manifesting in the ability to generate innovative ideas without additional learning (Franceschelli and Musolesi, 2023), indicating a degree of self-directed exploration and decision-making (Xi et al., 2023). Despite LLM-based agent’s application in diverse real-world scenarios such as software development (Li et al., 2023a; Qian et al., 2023), scientific research (Boiko et al., 2023), and system management (Liu et al., 2023), recent studies have highlighted issues of non-executable plans being generated without adequate human oversight (Ge et al., 2023a; Yuan et al., 2024). If these plans lack executability, the utility of LLM-based agents in fields that require high validity is compromised, and their unreliability erodes user trust. To address this challenge, we propose integrating natural language and precise automaton during the LLM-based agent’s planning process.
@@ 2.2 Controllable LLM Generation
To the best of our knowledge, there are few works on controllable LLM-based agents, and discussions primarily revolve around controllable LLM text generation. Controlling LLM text generation typically falls into two categories: hard constraints and soft constraints (Qin et al., 2022). Hard constraints involve constraints on the desired text length, designated keywords for the generated text (Takase and Okazaki, 2019; Carlsson et al., 2022), and constraint vocabulary space in decoding time (Hemmer et al., 2023; Geng et al., 2024, 2023), while soft constraints restrict the output based on specific semantics, such as sentiments or topics (Gu et al., 2022; Lu et al., 2022; Li et al., 2020). A recent work (Chen et al., 2024) attempts to integrate both constraints into a unified approach. However, text generation control has disadvantages and problems. First, text generation control could hardly be directly applied to closed-source LLMs due to their need to access the decoding probability (Geng et al., 2024). Second, controllable LLM-based agents prioritize the validity of plan and tool use over purely text generation. Recent studies on LLM-based agents have noted limited effectiveness when employing text constraint generation on LLM-based agents and have explored using LLMs as a parser to extract planning information from generated texts (Ge et al., 2023a; Yuan et al., 2024); but relying heavily on the parser LLM’s effectiveness may not yield a satisfactory plan valid rate. We propose an automaton-guided agent planning to ensure that the generated plans adhere 100% to the constraints.
@@ 3.1 Context-free Grammar
Many LLM agents operate tools that take multiple inputs and produce one output. For example, the tool “Visual Question Answering” receives an image and a text as the input and outputs a text. The plan involving such tools forms a tree structure, and such tree-structured plans require Context-Free Grammar (CFG) to express the rules.
The CFG consists of four components: terminals (symbols unable to be replaced), nonterminals (symbols subject to replacement), start symbol (a unique nonterminal, usually denoted as $S$), and productions (rules governing symbol substitution). The CFG production format follows:
@@
where $A$ is a single nonterminal, and $\alpha$ is a string comprising any combination of terminals and nonterminals. $A$ can be replaced by $\alpha$ in any situation.
$\begin{split}&S\rightarrow\varepsilon|aSb\end{split}$ (2)
Eq.(2) is a CFG example with two productions separated by “||||”, where $a,b$ are terminals, $S$ is the start symbol and the only nonterminal, and $\varepsilon$ stands for the empty string. From the start symbol, we can use these two productions to generate infinite strings, such as the empty string, $ab$, and $aabb$. Take the derivation of $aabb$ as an example. We first apply the second production twice ($S\rightarrow aSb\rightarrow aaSbb$) and then the first production once ($aaSbb\rightarrow aabb)$. The formal language, Context-Free Language (CFL), contains the set of strings of terminals, which can be derived from $S$ in a certain amount of CFG production steps.
Figure 2: A Pushdown Automaton (PDA) example.
Figure 3: An equivalent PDA to the CFG of the combination of Eq.(3) and Eq.(4). The lowercase letters $a$ to $f$ are $a_{i}$ to $f_{j}$ in Eq.(3). For example, the transit $(b,B;\varepsilon)$ implies the PDA removes the stack top $B$ and remains in state $q_{1}$ if the PDA is at state $q_{1}$, the stack top is $B$, and the next string letter is $b_{1}/b_{2}/b_{3}$.
@@ 3.2 Pushdown Automaton
Figure 2 shows how PDA works as an example. This automaton includes a stack alphabet ($\{S,Z\}$) with an initial stack symbol $Z$, a start state $q_{0}$, an accept state set $\{q_{2}\}$, and a transit function over states. The transit functions are in the form of triplets. For example, the edge $(a,Z;SZ)$ from $q_{0}$ to $q_{0}$ means that if the stack top is $Z$ and the next letter of the string is $a$, the PDA remains in $q_{0}$ state and the stack top changes from $Z$ to $SZ$ ($S$ is the stack top now).
@@ 4.1 Motivation and Challenge
Natural language is easily understandable by humans, but it may also lack precision in certain application scenarios. In contrast, formal language is defined in clear mathematical and machine readable forms, which could be hard to comprehend for humans, but it also possesses great precision. Thus, we aim to effectively integrate the benefits of both natural and formal languages through the pushdown automaton (PDA) for more controllable planning of LLM-based agents in scenarios requiring precise and valid plans.
Constructing our Formal-LLM framework presents non-trivial challenges. First, LLMs may struggle to directly comprehend and process formal language due to the limited CFG- or PDA-related corpus in pre-training data. Hence, natural language prompts are necessary to effectively describe the status of the automaton, bridging the gap between formal and natural language for LLM’s planning. Given that the automaton has the potential to generate an infinite number of plans, the second challenge arises because the automaton only ensures the validity and executability of the plans by adhering to defined constraints, but cannot guarantee the performance and optimality of the generated plan. Therefore, leveraging the solid natural language understanding ability of LLM and other methods, such as fine-tuning and backtracking mechanisms, becomes crucial for enhancing planning performance.
@@ 4.2 Formulating Constraint to Automaton
If the plans for the benchmark tasks generated by the LLM-based agent deviate from the expected data format, the execution performance degrades significantly or even results in errors. For instance, the Colorization tool operates on image data as input and output. Since image data is represented as a 3-D tensor, providing a 1-D tensor text string as input could lead to errors, due to incorrect data dimensions. Thus, the entire plan becomes invalid and cannot be executed. In such cases, natural language may struggle to articulate these constraints precisely and concisely, but formal language can express them straightforwardly.
@@
The CFG in Eq.(3) outlines the constraints on data formats with Polish prefix notation: $T$ for text and $I$ for image, capital letters $A$ to $F$ for the six types of tools, and lowercase letters such as $a_{i}\in A$ or $f_{j}\in F$ for specific tools in the order of Table 3.
$I\rightarrow AI|CT\qquad$ $T\rightarrow BI|DT|EIT|FTT$ (3)
$A\rightarrow a_{1}|a_{2}|a_{3}|a_{4}\qquad$ $B\rightarrow b_{1}|b_{2}|b_{3}$
$C\rightarrow c_{1}\qquad$ $D\rightarrow d_{1}|d_{2}|d_{3}|d_{4}|d_{5}$
$E\rightarrow e_{1}\qquad$ $F\rightarrow f_{1}$
@@
where $I\rightarrow AI|CT$ means $I$ can be replaced by $AI$ or $CT$, because $A$ is an image-to-image tool and $C$ is a text-to-image tool. As a result, when $A$ is applied on an image ($AI$) or $C$ is applied on a text ($CT$), the result is still an image. Considering a specific task, such as “Given blurry gray-scale images, how to return the object names in English step by step?”, we know that the task’s inputs are images and the final outputs should be texts. To formalize this, we add two constraints to Eq.(3):
$S\rightarrow T\qquad$ $I\rightarrow i$ (4)
@@
i.e., the CFG starts with the symbol $T$, implying the final text format, and uses the lowercase letter $i$ to represent the input image. The combination of Eq.(3) and Eq.(4) renders the data format constraints of the specific task. For instance, $e_{1}a_{1}ib_{1}i$ is an accepted word, representing a valid plan: “Utilize Image Classification ($b_{1}$) on the input image to obtain text and apply Colorization ($a_{1}$) on the input image to generate the output image. Then, employ Visual Question Answering ($e_{1}$) with the text and output image to derive the final text.” As mentioned in Section 3.2, the CFG can be equivalently converted into a PDA in Figure 3. These steps allow us to transform the constraints into automata for subsequent utilization.
@@
In certain scenarios, creating an automaton can be more straightforward than formulating a formal language. Section A.5.1 describes a real-life task for daily work planning. The constraints in this case revolve around time, allowing us to represent time as distinct states, as depicted in Figure 4. The automaton initiates the design of an activity that concludes at 20:00, and a valid plan is generated when the automaton reaches 10:00 while being in a state where three meals have been consumed. In this case, we design a PDA without a CFG.